13 A series LC circuit contains a 500 mHinductor a 25.0 mF capacitor, and a 13) 12 volt battery. What is the frequency of the electromagnetic oscillations?
Answer #1

Frequency of oscillation

f=\frac{1}{2\pi }\int \sqrt{\frac{1}{LC}}=\frac{1}{2\pi }\int \sqrt{\frac{1}{500*10^{-3}*25*10^{-3}}}

f=1.42hz